OpenStudy (greenpuppyz):

Divide mentally. 1.2 ÷ 3,000 = ? A. 0.4 B. 0.04 C. 0.004 D. 0.0004

OpenStudy (anonymous):

but hold ob im going to show you how

OpenStudy (er.mohd.amir):

1.2/3000=12/30000=12/(3*10000)=4/10000=.0004

OpenStudy (anonymous):

:000

OpenStudy (anonymous):

@alibaby is wrong.

OpenStudy (anonymous):

: 000 !!

OpenStudy (anonymous):

how?

OpenStudy (greenpuppyz):

its D i figured it out

OpenStudy (anonymous):

@GreenPuppyz when you are dividing by 1000, you have to move 3 decimal places to the left!

OpenStudy (lolacole12):

1.2/3000 =0.0004

OpenStudy (lolacole12):

yes it is D like i said gj @GreenPuppyz can i have a medal

OpenStudy (anonymous):

okay so its d?

OpenStudy (lolacole12):

yay

OpenStudy (anonymous):

my bad wow but it's 4 sorry @GreenPuppyz

OpenStudy (lolacole12):

i know its kinda hard

OpenStudy (anonymous):

0.0004

OpenStudy (lolacole12):

gj @alibaby and @GreenPuppyz or anyone need help make a question and call me

OpenStudy (anonymous):

c

OpenStudy (anonymous):

the answer is c

OpenStudy (lolacole12):

its not c

OpenStudy (anonymous):

it is

OpenStudy (greenpuppyz):

google said D

OpenStudy (anonymous):

(/O-O\) oh dear

OpenStudy (lolacole12):

not read the messages and yay soo

OpenStudy (anonymous):

u r a cheater because u searched it on the internet

OpenStudy (anonymous):

What

OpenStudy (anonymous):

what does u r a mean?

OpenStudy (anonymous):

you are a

OpenStudy (anonymous):

oh...

OpenStudy (lolacole12):

but he did to make sure we were right and we were so please do come in and do that if your not right and we tell you you say oh ok and you learn somthing you dont keep say that were wrong

OpenStudy (anonymous):

i am really sorry (/oo/)

OpenStudy (lolacole12):

Ok

OpenStudy (lolacole12):

its ok

OpenStudy (anonymous):

y r u a human calculator?

OpenStudy (lolacole12):

who me cause i help alot

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

u r kind and helpfull

OpenStudy (anonymous):

sooooo its D?

OpenStudy (lolacole12):

ok lol shure but you got a question other than y u a human calculater

OpenStudy (lolacole12):

@Gingerness yes lol

OpenStudy (anonymous):

so if i help a lot will i be one i dont think so because i started uesterday

OpenStudy (anonymous):

he got enough smartscore to where he could buy the title

OpenStudy (lolacole12):

well i gtg help others bye guys and maybye help 2 hours a day and you should get it or if you yay

OpenStudy (lolacole12):

get like the coins smart coins you can buy it

OpenStudy (anonymous):

thank u ever so much